# A Framework for Quantum Simulations of Energy-Loss and Hadronization in Non-Abelian Gauge Theories: SU(2) Lattice Gauge Theory in 1+1D

### v2

@@ V.1.3 Heavy-Quark Translation and Time Evolution
where $(\hat{\mathcal{Q}}_{n}^{(a)})^{2}$ corresponds to single-staggered-site contributions, while $\hat{\mathcal{Q}}_{n}^{(a)}\hat{\mathcal{Q}}_{m}^{(a)}$ corresponds to contributions from different staggered sites. Same-site interactions contain only $\hat{Z}\hat{Z}$ operators and can be implemented straightforwardly. Interactions between different staggered sites can be implemented efficiently using a GHZ transformation [Farrell:2022vyh]. [^11] The GHZ transformation $\mathcal{G}$ used here is to be distinguished from the GHZ transformation $G$ use earlier in the state preparation circuits. They facilitate the same goal of implementing a target operator by transforming the basis of efficient circuits in the computational basis. Since the target operators in state preparation and time evolution have different forms, the GHZ transformations are slightly different The operator order from left to right correspond to circuit wires from top to bottom.
@@
This does not include truncations of the chromo-electric contributions made possible by confinement [Farrell:2024fit], as the chromo-electric Hamiltonian is symmetrically implemented, and interactions act on a maximum of two spatial sites. We perform a second order Trotterization, where each step of the time evolution is
@@ V.3 Results from Quantum Simulations
A suite of state preparation and time evolution circuits, along with the associated error-mitigation circuits, that have been detailed in previous sections, were implemented with IBM’s qiskit [qiskit] and executed on IBM’s quantum computer ibm_pittsburgh. The results of these simulations were analyzed classically with standard statistical techniques using python, julia and Mathematica. We focus on two observables related to heavy-quark motion: charge distribution and energy loss.
@@ V.3.1 Charge Distribution
An initial state of the $L=3$ ($n_{\rm qubits}=18$) system using Hamiltonian parameters $m_{q}=0.1$ and $g=1.0$ was prepared with a heavy quark residing on the $x=0$ spatial site. The heavy quark is then moved to the $x=1$ spatial site and the system is time evolved forward to $t=2.0$ using a single second-order Trotter step. Several error mitigation methods were applied to compensate for the noisy hardware. Pauli twirling [Wallman:2015uzh] was performed around the entangling gates, with a total of 32 distinct twirled circuits for each of the two independent production runs. Dynamical decoupling [Viola:1998jx, Ezzell:2022uat] and measurement twirling [Berg:2020ibi] were also used. To remove the remaining bias due to incoherence of the device, we apply operator-decoherence renormalization (ODR) [Urbanek:2021oej, ARahman:2022tkr, Farrell:2023fgd]. For that, we ran two types of circuits: the “physics” circuit that implements the process of interest, and two “mitigation” circuits. The mitigation circuits are set up such that they share the same structure as the physics circuits, but are easy to simulate classically in that they only contain Clifford circuits. In the state preparation, the non-Clifford angles in the rotation gates [^13] The gates $R_{\pm}^{XY}(\theta)$ and $R_{\pm}^{XX}(\theta)$ with $\theta=\pm\frac{\pi}{2}$ in Fig. 13 do not get modified for the mitigation circuits. are set to $\theta_{\rm mit}=0,\pi/2$. In the time evolution, the mitigation circuits evolve the system forward then backward in time such that there is no change in an ideal simulation. [^14] For one step of second-order Trotterized time evolution, the mitigation circuit becomes $e^{i\frac{t}{2}\hat{H}_{k1}}e^{i\frac{t}{2}\hat{H}_{k0}}e^{-i0\hat{H}_{m}}e^{-i0\hat{H}_{g}}e^{-i\frac{t}{2}\hat{H}_{k0}}e^{-i\frac{t}{2}\hat{H}_{k1}}\ .$ (65) With these quantities, the rescaling factor to compensate for the decoherence can be computed:
@@
The quantum circuits were transpiled to run on IBM hardware. For each twirl of each type of circuit, 4096 measurements (shots) of $\langle\hat{Z}_{j}\rangle$ were performed in the computational basis. For each qubit, we obtain three results (one from the physics circuit, and two from the mitigation circuits), each for a total of 64 twirls. For the mitigation circuits, $\langle\hat{Z}_{j}\rangle^{\rm mit}_{\rm pred}$ is expected to take three values, $\{-1,0,1\}$. The reason we chose to run two different mitigation circuits, is that with each of the mitigation circuits, some of the observables can take the value zero, incompatible with ODR. With the two choices of $\theta_{\rm mit}$ selected, we have at least one non-zero value for $\langle\hat{Z}_{j}\rangle^{\rm mit}_{\rm pred}$. Measurements with $|\langle\hat{Z}_{j}\rangle^{\rm mit}_{\rm meas}|<\lambda^{\rm mit}_{\rm cut}$ and $|\langle\hat{Z}_{j}\rangle^{\rm phys}_{\rm meas}|<\lambda^{\rm phys}_{\rm cut}$, are filtered out [Farrell:2024fit, Farrell:2025nkx]. Additionally, the inherent SU(2) symmetry of the system, which requires that pairs of qubits within a staggered site to have the same value of $\langle\hat{Z}_{j}\rangle$, was used to combined results for the $r,g$ qubits. The median and (re-scaled) median-average-deviation (MAD) were used as robust estimators of the predicted central value of $\langle\hat{Z}_{j}\rangle$ for each qubit, along with its $68\%$ confidence interval. The value of $\lambda^{\rm mit}_{\rm cut}$ and $\lambda^{\rm phys}_{\rm cut}$ was set at $0.005$, data below which is considered unrecoverable for a circuit with $\sim 400$ two-qubit-gate depth. The results of this analysis are compiled in Table LABEL:tab:IBManalysis, and displayed in Fig. 19.
Figure 19: Results obtained from ibm_pittsburgh for $\langle\hat{Z}_{j}\rangle$ in the $L=3$ system. At $t=0$, the heavy quark is instantaneously moved from $x=0$ to $x=1$, then evolved forward in time to $t=2.0$ using one step of the second-order Trotterized time-evolution operator. The SU(2) symmetry has been used to combine results for each red and green quark. The left panel shows the expectation values obtained using bootstrap resampling with the median and MAD. The points with error bars correspond to the raw and mitigated results obtained from ibm_pittsburgh, while the gray bar outlines are the expected results obtained using classical computing. The right panel shows histograms of results for $\langle\hat{Z}_{j}\rangle$ for each pair of qubits. Each color corresponds to a different staggered site.
@@ V.3.2 Energy Loss
The energy transfer to the system due to the heavy-quark translating from $x=0$ to $x=1$ was determined by running three different physics circuits, along with corresponding mitigation circuits, [^16] The mitigation circuits for the operators in $\hat{P}_{1}$ and $\hat{P}_{2}$ have the non-Clifford rotation angles set to 0, while no mitigation circuits were run for $\hat{P}_{3}$. This is because, for generating the $\hat{P}_{3}$ mitigation circuits, setting non-Clifford rotation angles to 0 or $\pi/2$ both result in 0 for noiseless predictions of observables, incompatible with ODR. on the quantum computer. As the last operator in the state preparation circuit does not change the measurements, a shallower circuit, with a CZ depth of 115 compared to 138, is used to measure the diagonal Pauli strings defined in Eq. (68).
In addition to the error mitigation methods applied in the charge distribution measurements, zero-noise extrapolation (ZNE) [ZNE1, ZNE2, PhysRevLett.120.210501, Klco:2018kyo, ZNE3] is employed through qiskit Runtime options [qiskit]. Noise factors-$(1,2,3)$ are executed on the device, where the different noise factors indicate the increase in circuit depth due to additional CZ gates (e.g., factor 1 is the original circuit, and factor 3 has three times the CZ count of the original circuit). In post-processing, and before performing ZNE of the raw data, we perform ODR for each noise level using the 32 sets of Pauli-twirled physical and mitigation circuits. These distributions are then linearly extrapolated in the noise-factor to the point of zero noise, providing an error-mitigated estimate of the quantity of interest. Computing the three contributions to the energy difference given in Eq. (68), both with and without ZNE for the sake of comparison, gives,
$\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}_{t<0}$ $=0.3755(11)-0.7275(32)-0.0326(05)\ =\ -0.3846(34)\ \ ({\rm ODR})\,$ (69)
$\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}_{t<0}$ $=0.3393(26)-0.8270(77)-0.0431(10)\ =\ -0.5308(89)\ \ ({\rm ODR+ZNE})\ ,$ (70)
which are to be compared with the exact values, $\langle\Delta\hat{H}_{g}\rangle^{\rm exact}_{t<0}=0.3314-0.7951-0.0421=-0.5058$, obtained from a noiseless simulation of the quantum circuits. Combined with $\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}_{0<t<1}=0$, this result implies an energy loss of $\delta\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}=-0.5301(48)$ due to the translation of the heavy quark from $x=0$ to $x=1$. Further details about this determination can be found in App. LABEL:app:IBMruns. It is important to note the importance of combining ODR and ZNE in obtaining a result from ibm_pittsburgh that is close to the expected value. This is, in part, due to the modest cancellations among contributions to the sum, that magnify the impact of the individual errors.
@@ VI Summary and Conclusions
With all the elements in place to perform measurements of energy loss in non-Abelian systems in 1+1D, there are a number of near-term directions to pursue. One direction is to perform simulations on larger lattices, with the heavy quark initially located far from the boundaries. This will permit a more robust implementation of domain decomposition for initial state preparation. Further, it will allow for estimates of energy loss with reduced systematic uncertainties, and changing the Hamiltonian parameters to decrease the lattice spacing. Hadrons created by the motion of the heavy quark will have more time to propagate before encountering the boundaries and reflecting back to the heavy quark, providing an extended time interval to study the fundamental mechanisms without interference. A second direction is to simulate energy loss and hadronization in the analogous SU(3) lattice gauge theory. Heavy quarks can now be included in existing frameworks of 1+1D SU(3) simulations [Ciavarella:2021nmj, Farrell:2022wyt, Farrell:2022vyh, Atas:2022dqm, Chernyshev:2025lil]. Such simulations with one flavor of light quarks require 9 qubits per spatial site (rgbrgb for the heavy and light quarks, and $\bar{r}\bar{g}\bar{b}$ for the light anti-quarks), and two-flavor simulations require 15 qubits per spatial site. A third direction, which is less trivial and may require the development of additional techniques, is the extension to simulations in 2+1D and 3+1D. Significant progress is being made in evolving quantum simulations of lattice field theories into 2+1D, with explorations in 3+1D in limited cases [Raychowdhury:2018osk, Ciavarella:2021nmj, Zache:2023dko, Muller:2023nnk, Kavaki:2024ijd, Turro:2024pxu, Gustafson:2023aai, Kadam:2024ifg, Fontana:2024rux, Grabowska:2024emw, Ciavarella:2024cyt, Halimeh:2024bth, Ciavarella:2024fzw, Balaji:2025afl, Illa:2025dou, Illa:2025njz, Jiang:2025ufg, Yao:2025uxz]. Typically, these involve simulating small volumes of the Kogut-Susskind Hamiltonian in the electric basis, with a significantly truncated gauge space, and without quantification of systematic errors. However, full simulations with both quarks and gluons remain to be performed. Such simulations are a prerequisite for qualitative comparisons with experiment and observation, and for making robust predictions in regimes where experiment is not possible.
@@ Acknowledgements.
We would like to thank Roland Farrell, Henry Froland, Nikita Zemlevskiy and Roland de Putter for helpful discussions. This work was supported, in part, by U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) [^17] https://iqus.uw.edu under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science, [^18] https://science.osti.gov/np/Research/Quantum-Information-Science by the Quantum Science Center (QSC) [^19] https://www.qscience.org which is a National Quantum Information Science Research Center of the U.S. Department of Energy, and by PNNL’s Quantum Algorithms and Architecture for Domain Science (QuAADS) Laboratory Directed Research and Development (LDRD) Initiative. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830. It was also supported, in part, by the Department of Physics [^20] https://phys.washington.edu and the College of Arts and Sciences [^21] https://www.artsci.washington.edu at the University of Washington. This research used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington. [^22] https://itconnect.uw.edu/research/hpc We have made extensive use of Wolfram Mathematica [Mathematica], python [python3, Hunter:2007], julia [Julia-2017], jupyter notebooks [PER-GRA:2007] in the Conda environment [anaconda], and IBM’s quantum programming environment qiskit [qiskit]. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award NP-ERCAP0032083.

### v3

@@ V.1.3 Heavy-Quark Translation and Time Evolution
where $(\hat{\mathcal{Q}}_{n}^{(a)})^{2}$ corresponds to single-staggered-site contributions, while $\hat{\mathcal{Q}}_{n}^{(a)}\hat{\mathcal{Q}}_{m}^{(a)}$ corresponds to contributions from different staggered sites. Same-site interactions contain only $\hat{Z}\hat{Z}$ operators and can be implemented straightforwardly. Interactions between different staggered sites can be implemented efficiently using a GHZ transformation [42]. [^11] The GHZ transformation $\mathcal{G}$ used here is to be distinguished from the GHZ transformation $G$ use earlier in the state preparation circuits. They facilitate the same goal of implementing a target operator by transforming the basis of efficient circuits in the computational basis. Since the target operators in state preparation and time evolution have different forms, the GHZ transformations are slightly different The operator order from left to right correspond to circuit wires from top to bottom.
@@
This does not include truncations of the chromo-electric contributions made possible by confinement [53], as the chromo-electric Hamiltonian is symmetrically implemented, and interactions act on a maximum of two spatial sites. We perform a second order Trotterization, where each step of the time evolution is
@@ V.3 Results from Quantum Simulations
A suite of state preparation and time evolution circuits, along with the associated error-mitigation circuits, that have been detailed in previous sections, were implemented with IBM’s qiskit [154] and executed on IBM’s quantum computer ibm_pittsburgh. The results of these simulations were analyzed classically with standard statistical techniques using python, julia and Mathematica. We focus on two observables related to heavy-quark motion: charge distribution and energy loss.
@@ V.3.1 Charge Distribution
An initial state of the $L=3$ ($n_{\rm qubits}=18$) system using Hamiltonian parameters $m_{q}=0.1$ and $g=1.0$ was prepared with a heavy quark residing on the $x=0$ spatial site. The heavy quark is then moved to the $x=1$ spatial site and the system is time evolved forward to $t=1.0$ using a single second-order Trotter step. Several error mitigation methods were applied to compensate for the noisy hardware. Pauli twirling [155] was performed around the entangling gates, with 64 distinct twirled circuits for the production runs. Dynamical decoupling [156, 157] and measurement mitigation [158] were applied through the qiskit [154] Estimator primitive. To remove the remaining bias due to incoherence of the device, we apply operator-decoherence renormalization (ODR) [159, 38, 48]. For that, we ran two types of circuits: the “physics” circuit that implements the process of interest, and two “mitigation” circuits. The mitigation circuits are set up such that they share the same structure as the physics circuits, but are easy to simulate classically in that they only contain Clifford circuits. In the state preparation, the non-Clifford angles in the rotation gates [^13] The gates $R_{\pm}^{XY}(\theta)$ and $R_{\pm}^{XX}(\theta)$ with $\theta=\pm\frac{\pi}{2}$ in Fig. 13 do not get modified for the mitigation circuits. are set to $\theta_{\rm mit}=0,\pi/2$. In the time evolution, the mitigation circuits evolve the system forward then backward in time such that there is no change in an ideal simulation. [^14] For one step of second-order Trotterized time evolution, the mitigation circuit becomes $e^{i\frac{t}{2}\hat{H}_{k1}}e^{i\frac{t}{2}\hat{H}_{k0}}e^{-i0\hat{H}_{m}}e^{-i0\hat{H}_{g}}e^{-i\frac{t}{2}\hat{H}_{k0}}e^{-i\frac{t}{2}\hat{H}_{k1}}\ .$ (65) With these quantities, the rescaling factor to compensate for the decoherence can be computed:
@@
The quantum circuits were transpiled to run on IBM hardware. For each twirl of each type of circuit, 4096 measurements (shots) of $\langle\hat{Z}_{j}\rangle$ were performed in the computational basis. For each qubit, we obtain three results (one from the physics circuit, and two from the mitigation circuits), each for a total of 64 twirls. For the mitigation circuits, $\langle\hat{Z}_{j}\rangle^{\rm mit}_{\rm pred}$ is expected to take three values, $\{-1,0,1\}$. The reason we chose to run two different mitigation circuits, is that with each of the mitigation circuits, some of the observables can take the value zero, incompatible with ODR. With the two choices of $\theta_{\rm mit}$ selected, we have at least one non-zero value for $\langle\hat{Z}_{j}\rangle^{\rm mit}_{\rm pred}$. Measurements with $|\langle\hat{Z}_{j}\rangle^{\rm mit}_{\rm meas}|<\lambda^{\rm mit}_{\rm cut}$ and $|\langle\hat{Z}_{j}\rangle^{\rm phys}_{\rm meas}|<\lambda^{\rm phys}_{\rm cut}$, are filtered out [53, 66]. Additionally, the inherent SU(2) symmetry of the system, which requires that pairs of qubits within a staggered site to have the same value of $\langle\hat{Z}_{j}\rangle$, was used to combined results for the $r,g$ qubits. The median and (re-scaled) median-average-deviation (MAD) were used as robust estimators of the predicted central value of $\langle\hat{Z}_{j}\rangle$ for each qubit, along with its $68\%$ confidence interval. The value of $\lambda^{\rm mit}_{\rm cut}$ and $\lambda^{\rm phys}_{\rm cut}$ was set at $0.005$, data below which is considered unrecoverable for a circuit with $\sim 400$ two-qubit-gate depth. The results of this analysis are compiled in Table LABEL:tab:IBManalysis, and displayed in Fig. 19.
Figure 19: Results obtained from ibm_pittsburgh for $\langle\hat{Z}_{j}\rangle$ in the $L=3$ system. At $t=0$, the heavy quark is instantaneously moved from $x=0$ to $x=1$, then evolved forward in time to $t=1.0$ using one step of the second-order Trotterized time-evolution operator. The SU(2) symmetry has been used to combine results for each red and green quark. The left panel shows the expectation values of $\langle\hat{Z}_{j}\rangle$ before the heavy quark is moved; the right panel shows the values at $t=1$. The results are obtained using bootstrap resampling with the median and MAD. The points with error bars correspond to the raw (orange circle) and mitigated (blue line) results obtained from ibm_pittsburgh, while the gray bar outlines are the expected results obtained using classical computing.
@@ V.3.2 Energy Loss
The energy transfer to the system due to the heavy-quark translating from $x=0$ to $x=1$ was determined by running three different physics circuits, along with corresponding mitigation circuits, [^16] The mitigation circuits for the operators in $\hat{P}_{1}$ and $\hat{P}_{2}$ have the non-Clifford rotation angles set to 0, while no mitigation circuits were run for $\hat{P}_{3}$. This is because, for generating the $\hat{P}_{3}$ mitigation circuits, setting non-Clifford rotation angles to 0 or $\pi/2$ both result in 0 for noiseless predictions of observables, incompatible with ODR. on the quantum computer. As the last operator in the state preparation circuit does not change the measurements, a shallower circuit, with a CZ depth of 109, is used to measure the diagonal Pauli strings defined in Eq. (68).
In addition to the error mitigation methods applied in the charge distribution measurements, zero-noise extrapolation (ZNE) [160, 161, 162, 18, 163] is employed through qiskit Runtime options [154]. Noise factors-$(1,2,3)$ are executed on the device, where the different noise factors indicate the increase in circuit depth due to additional CZ gates (e.g., factor 1 is the original circuit, and factor 3 has three times the CZ count of the original circuit). In post-processing, and before performing ZNE of the raw data, we perform ODR for each noise level using the 32 sets of Pauli-twirled physical and mitigation circuits. These distributions are then linearly extrapolated in the noise-factor to the point of zero noise, providing an error-mitigated estimate of the quantity of interest. Computing the three contributions to the energy difference given in Eq. (68), both with and without ZNE for the sake of comparison, gives,
$\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}_{t<0}$ $=0.3715(16)-0.7243(30)-0.0314(04)\ =\ -0.3841(30)\ \ ({\rm ODR})\,$ (69)
$\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}_{t<0}$ $=0.3491(19)-0.7538(73)-0.0390(07)\ =\ -0.4437(72)\ \ ({\rm ODR+ZNE})\ ,$ (70)
which are to be compared with the exact values, $\langle\Delta\hat{H}_{g}\rangle^{\rm exact}_{t<0}=0.3314-0.7951-0.0421=-0.5058$, obtained from a noiseless simulation of the quantum circuits. Combined with $\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}_{0<t<1}=0$, this result implies an energy loss of $\delta\langle\Delta\hat{H}_{g}\rangle^{\rm IBM}=+0.4437(72)$ due to the translation of the heavy quark from $x=0$ to $x=1$. Further details about this determination can be found in App. LABEL:app:IBMruns. It is important to note the importance of combining ODR and ZNE in obtaining a result from ibm_pittsburgh that is close to the expected value. This is, in part, due to the modest cancellations among contributions to the sum, that magnify the impact of the individual errors.
@@ VI Summary and Conclusions
With all the elements in place to perform measurements of energy loss in non-Abelian systems in 1+1D, there are a number of near-term directions to pursue. One direction is to perform simulations on larger lattices, with the heavy quark initially located far from the boundaries. This will permit a more robust implementation of domain decomposition for initial state preparation. Further, it will allow for estimates of energy loss with reduced systematic uncertainties, and changing the Hamiltonian parameters to decrease the lattice spacing. Hadrons created by the motion of the heavy quark will have more time to propagate before encountering the boundaries and reflecting back to the heavy quark, providing an extended time interval to study the fundamental mechanisms without interference. A second direction is to simulate energy loss and hadronization in the analogous SU(3) lattice gauge theory. Heavy quarks can now be included in existing frameworks of 1+1D SU(3) simulations [24, 40, 42, 41, 69]. Such simulations with one flavor of light quarks require 9 qubits per spatial site (rgbrgb for the heavy and light quarks, and $\bar{r}\bar{g}\bar{b}$ for the light anti-quarks), and two-flavor simulations require 15 qubits per spatial site. A third direction, which is less trivial and may require the development of additional techniques, is the extension to simulations in 2+1D and 3+1D. Significant progress is being made in evolving quantum simulations of lattice field theories into 2+1D, with explorations in 3+1D in limited cases [164, 24, 165, 166, 54, 56, 167, 168, 169, 170, 171, 172, 57, 173, 174, 175, 114, 176]. Typically, these involve simulating small volumes of the Kogut-Susskind Hamiltonian in the electric basis, with a significantly truncated gauge space, and without quantification of systematic errors. However, full simulations with both quarks and gluons remain to be performed. Such simulations are a prerequisite for qualitative comparisons with experiment and observation, and for making robust predictions in regimes where experiment is not possible.
@@ Acknowledgements.
We would like to thank Roland Farrell, Henry Froland, Nikita Zemlevskiy and Roland de Putter for helpful discussions. This work was supported, in part, by U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) [^17] https://iqus.uw.edu under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science, [^18] https://science.osti.gov/np/Research/Quantum-Information-Science by the Quantum Science Center (QSC) [^19] https://www.qscience.org which is a National Quantum Information Science Research Center of the U.S. Department of Energy, and by PNNL’s Quantum Algorithms and Architecture for Domain Science (QuAADS) Laboratory Directed Research and Development (LDRD) Initiative. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830. It was also supported, in part, by the Department of Physics [^20] https://phys.washington.edu and the College of Arts and Sciences [^21] https://www.artsci.washington.edu at the University of Washington. This research used resources of the Oak Ridge Leadership Computing Facility (OLCF), which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington. [^22] https://itconnect.uw.edu/research/hpc We have made extensive use of Wolfram Mathematica [177], python [178, 179], julia [180], jupyter notebooks [181] in the Conda environment [182], and IBM’s quantum programming environment qiskit [154]. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award NP-ERCAP0032083.